OpenStudy (anonymous):

Evaluate f(x) = x 2 + 1 for f(-1). a.0 b.undefined c.2

OpenStudy (snuggielad):

Do this the same way you did it before. X=1...put -1 in for x in that problem

OpenStudy (snuggielad):

(-1)*2+1

OpenStudy (snuggielad):

or also written it is 2(-1) + 1

hero (hero):

f(x) = x^2 + 1 f(-1)= (-1)^2 + 1 = 1 + 1 = ?

OpenStudy (snuggielad):

Hold on? Is that a square root? Be sure to write that out next time

OpenStudy (snuggielad):

Well done @Hero

hero (hero):

Its \(f(x) = x^2 + 1\) Sometimes when students copy and paste, they forget to add the caret.

OpenStudy (anonymous):

ok so the answer is 2 thanks a lot

OpenStudy (snuggielad):

Yes, I should have picked up on that.

OpenStudy (anonymous):

can you guys help me with one more its Evaluate f(x) = x^ 2 + 1 for f(10). a.2√2 b.101 c.21

hero (hero):

f(x) = x^2 + 1 f(10) = (10)^2 + 1 = 100 + 1 = ?

OpenStudy (anonymous):

what about this one? If f(x) = x 2 + 1 and g(x) = 3x + 1, find [f(2)-g(1)] 2. a.1 b.2 c.9

hero (hero):

I thought you said you only had one more....You tricked me :)

OpenStudy (anonymous):

sorry...lol this is the last one.

hero (hero):

f(2) - g(1) = (2)^2 + 1 -(3(1) + 1) = 4 + 1 - 3 - 1 = 5 - 4 = ?

OpenStudy (anonymous):

thanks